OpenStudy (anonymous):

Solve. Square root of x+6=2-square root of x+2 I know you square both sides and get x+6=(2-square root x+2)(2-square root x+2) Help please! I'm stuck the answer is x= -2 But I don't know all the steps

OpenStudy (anonymous):

is it \[x+6=2-\sqrt{x+2}\]?

OpenStudy (anonymous):

if so first step is subtract 2 to get \[x+4=\sqrt{x+2}\] then square

OpenStudy (anonymous):

oh okay & then you... square both sides?

OpenStudy (anonymous):

wait no the x+6 is a square root

OpenStudy (anonymous):

so you can't subtract the 2 right? It'd be x+6=4-2squareroot x+2 - 2squareroot x+2+x+2

OpenStudy (anonymous):

well then you have to square both sides twice

OpenStudy (anonymous):

yes after doing that, I got what I just sent you is that right?

OpenStudy (anonymous):

\[x+6=4-4\sqrt{x+2}+x+2\]

OpenStudy (anonymous):

that is step one yes?

OpenStudy (anonymous):

Yes and then I would add the 2 to the 4 so then it would equal x+6=6+x-4squareroot x+2 right?

OpenStudy (anonymous):

now we get \[0=-4\sqrt{x+2}\] making \[\sqrt{x+2}=0\] or \[x+2=0\] \[x=-2\]

OpenStudy (anonymous):

you have an extra 2 hanging out at the end i think. 4+2 = 6

OpenStudy (anonymous):

Oh okay I get it! :D

OpenStudy (anonymous):

you wrote Yes and then I would add the 2 to the 4 so then it would equal x+6=6+x-4squareroot x+2 right? but i think you just get x+6=6+x-4 root(x+2)

OpenStudy (anonymous):

so answer is 0 and check is easy

OpenStudy (anonymous):

I have more if you have time? I have a test tomorrow :(

OpenStudy (anonymous):

yes I see where I messed up thank you

OpenStudy (anonymous):

\[\sqrt{2x+2}-\sqrt{x-3}=2\]

OpenStudy (anonymous):

I know you take x square root x-3 to the other side first

OpenStudy (anonymous):

so then it would be\[\sqrt{2x+2}^{2}=(2+\sqrt{x-3)^{2}}\]